# Corona rigidity

### v3

@@ 11.0.2. The role of MA and $\mathrm{OCA^{\#}}$
MA is not needed in order to prove that all automorphisms of the Calkin algebra are inner, because of the isometry trick (see Lemma 7.14). It is not known whether $\mathrm{OCA_{T}}$ suffices to prove that all automorphisms of $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ are trivial, but one does not need the full power of $\mathrm{MA}$. A variant of $\mathrm{OCA_{T}}$ called $\mathrm{OCA^{\#}}$ has been isolated in [32] and shown to follow from $\mathrm{OCA_{T}}$ and $\mathrm{MA_{\sigma-linked}}$. This axiom does not have the elegance of $\mathrm{OCA_{T}}$, but it is tailor-made for rigidity proofs and it alone provides a reasonably simple proof that all automorphisms of $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ are trivial. (In addition, the proof that $\mathrm{OCA_{T}}$ and $\mathrm{MA_{\sigma-linked}}$ together imply $\mathrm{OCA^{\#}}$ is very straightforward.) It is not known whether $\mathrm{OCA_{T}}$ implies that all automorphisms of $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ are trivial, or whether it implies $\mathrm{OCA^{\#}}$.
@@ 11.0.4. A ‘commutative’ version of the BDF question
for $n\in{\mathbb{N}}$ and $A\in\mathcal{P}({\mathbb{N}})$. Is there an automorphism $\Phi$ of $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ such that $1.\Phi([A])=\Phi(-1.[A])$ for all $A$? [^44] It should be noted that a positive answer to this question would not be a progress towards solving Question 4.16. This is because Truss’s theorem (see Example 2.8) implies that if an automorphism of the Calkin algebra sends the atomic masa $\ell_{\infty}({\mathbb{N}})/c_{0}$ to itself then its restriction to the normalizer of the atomic masa is implemented by a unitary; see [59, Notes to §17.9]. Such $\Phi$ cannot be trivial. Numerous partial positive results on this problem have been obtained in [77] and [153]. This and related problems were tackled from the dynamic side in [16] and [15]. A positive answer to this question was announced in [17].
@@ 11.0.6. Rigidity for ultrapowers
The abstract and general setup outlined in §2 is an attempt to capture the notion of a definable (Borel) quotient structure whose properties (automorphism group in particular) are sensitive to the choice of additional set-theoretic axioms. Needless to say, some quotient structures of general interest are not covered by our setup—for example, ultraproducts (see also §11.0.7). Throughout this survey, we considered only structures ${\mathcal{M}}/E$ for a Borel ${\mathcal{M}}$ and a Borel $E$. An important instance of the rigidity question is obtained by relaxing the Borelness requirement on $E$, is the rigidity question for ultrapowers. For a given Borel structure ${\mathcal{M}}$ (belonging to an axiomatizable category, so that the ultrapowers are defined) consider the Borel structure ${\mathcal{M}}^{\mathbb{N}}$. If $\mathcal{U}$ is an ultrafilter on ${\mathbb{N}}$, then the ultrapower ${\mathcal{M}}_{\mathcal{U}}$ of ${\mathcal{M}}$ associated with $\mathcal{U}$ is of the form ${\mathcal{M}}^{\mathbb{N}}/E_{\mathcal{U}}$ for an appropriately defined $E_{\mathcal{U}}$. The standard saturation arguments [^45] We assume that the ultrafilters $\mathcal{U}$ and $\mathcal{V}$ are nonprincipal.6.1) implies that, under $\mathrm{CH}$, ${\mathcal{M}}_{\mathcal{U}}$ has $2^{\aleph_{1}}$ automorphisms and that ${\mathcal{M}}_{\mathcal{U}}$ and ${\mathcal{N}}_{\mathcal{V}}$ are isomorphic if and only if ${\mathcal{M}}$ and ${\mathcal{N}}$ are elementarily equivalent. Whether this conclusion is true without the $\mathrm{CH}$ depends on whether the theory of ${\mathcal{M}}$ is stable or not (see [65] and [69]).
@@ 11.0.9. Maximal rigidity
In Example 2.8 we saw that another quotient group, $S_{\infty}/H$, (where $H$ denotes the normal subgroup of all finitely supported permutations of ${\mathbb{N}}$) has the property that (provably in $\mathrm{ZFC}$) each of its automorphisms is given as conjugation by an almost permutation, and in particular it has a continuous lifting. One consequence of this strong rigidity result is the following. Fix a pair $f$, $g$ in $S_{\infty}$, and let ${\mathcal{M}}=(S_{\infty},f)$ and ${\mathcal{N}}=(S_{\infty},g)$. Then ${\mathcal{M}}/H$ and ${\mathcal{N}}/H$ are isomorphic if and only if $f$ and $g$ are conjugate (by an almost permutation) in the quotient, and in particular the question of the isomorphism of these structures is independent from $\mathrm{ZFC}$.
@@
What other structures (or categories) have this rigid behaviour, and is it possible to give a general characterization of such structures?
@@ 12.1.2. BDF
The Brown–Douglas–Filmore question (§4.3), whether there exists an automorphism of the Calkin algebra that sends the unilateral shift to its adjoint, can be considered within the framework of (12.1). Considering ${\mathcal{M}}=(\mathcal{B}(H),s)$ and ${\mathcal{N}}=(\mathcal{B}(H),s^{*})$ as structures in the language of $\mathrm{C}^{*}$-algebras expanded by a constant interpreted by the shift or by its adjoint, the question is whether ${\mathcal{M}}/\mathcal{K}(H)$ is isomorphic to ${\mathcal{N}}/\mathcal{K}(H)$. The lamentable status of this question can be summarized in two remarks. First, it is not even known whether these two structures are elementarily equivalent. Second, even if they were, the Calkin algebra—and therefore each of these two structures—fails even to be countably homogeneous (see §6.3). Nevertheless, Woodin’s $\Sigma^{2}_{1}$ absoluteness theorem suggests that assuming the Continuum Hypothesis may help. In addition, forcing $\mathrm{CH}$ without adding reals preserves $\Sigma^{2}_{1}$ statements and therefore assuming $\mathrm{CH}$ does no harm in this context.
@@ 12.1.3. Non-isomorphism
In order to avoid metamathematical detours, instead of question (12.2) we consider a question that is, given the current state of the art in set theory, easier to analyze: When does /≇${\mathcal{M}}/{\mathcal{I}}\not\cong{\mathcal{M}}/{\mathcal{J}}$ hold in a forcing extension?

### v4

@@ 11.0.2. The role of MA, $\mathrm{OCA^{\#}}$, and Biba’s trick
MA is not needed in order to prove that all automorphisms of the Calkin algebra are inner, because of the isometry trick (see Lemma 7.14). In [32, Theorem 1] it was finally proved that $\mathrm{OCA_{T}}$ alone imples all automorphisms of $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ are trivial. The proof filters through a variant of $\mathrm{OCA_{T}}$ called $\mathrm{OCA^{\#}}$ that follows from $\mathrm{OCA_{T}}$ by a modification of Moore’s proof that $\mathrm{OCA_{T}}$ implies $\mathrm{OCA_{\infty}}$ ([120], [32, Theorem 3.3]). This axiom does not have the elegance of $\mathrm{OCA_{T}}$, but its complexity is generously compensated by its utility in analysing automorphisms of quotient structures, via the so-called Biba’s trick. It was used in [64] to prove that $\mathrm{OCA_{T}}$ and $\mathrm{MA_{\sigma-linked}}$ together imply the conclusion of the OCA Lifting Theorem (Theorem 7.8) and to show that all isomorphisms between quotients over countably $80$-determined ideals on ${\mathbb{N}}$ are topologically trivial, partially answering Question 7.11 from an earlier version of this paper.
@@ 11.0.4. A ‘commutative’ version of the BDF question
for $n\in{\mathbb{N}}$ and $A\in\mathcal{P}({\mathbb{N}})$. Is there an automorphism $\Phi$ of $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ such that $1.\Phi([A])=\Phi(-1.[A])$ for all $A$? [^45] It should be noted that a positive answer to this question would not be a progress towards solving Question 4.16. This is because Truss’s theorem (see Example 2.8) implies that if an automorphism of the Calkin algebra sends the atomic masa $\ell_{\infty}({\mathbb{N}})/c_{0}$ to itself then its restriction to the normaliser of the atomic masa is implemented by a unitary; see [61, Notes to §17.9]. Such $\Phi$ cannot be trivial. Numerous partial positive results on this problem have been obtained in [81] and [160]. This and related problems were tackled from the dynamic side in [16] and [15]. A positive answer to this question was announced in [17].
@@ 11.0.6. Rigidity for ultrapowers
The abstract and general setup outlined in §2 is an attempt to capture the notion of a definable (Borel) quotient structure whose properties (automorphism group in particular) are sensitive to the choice of additional set-theoretic axioms. Needless to say, some quotient structures of general interest are not covered by our setup—for example, ultraproducts (see also §11.0.7). Throughout this survey, we considered only structures ${\mathcal{M}}/E$ for a Borel ${\mathcal{M}}$ and a Borel $E$. An important instance of the rigidity question is obtained by relaxing the Borelness requirement on $E$, is the rigidity question for ultrapowers. For a given Borel structure ${\mathcal{M}}$ (belonging to an axiomatisable category, so that the ultrapowers are defined) consider the Borel structure ${\mathcal{M}}^{\mathbb{N}}$. If $\mathcal{U}$ is an ultrafilter on ${\mathbb{N}}$, then the ultrapower ${\mathcal{M}}_{\mathcal{U}}$ of ${\mathcal{M}}$ associated with $\mathcal{U}$ is of the form ${\mathcal{M}}^{\mathbb{N}}/E_{\mathcal{U}}$ for an appropriately defined $E_{\mathcal{U}}$. The standard saturation arguments [^46] We assume that the ultrafilters $\mathcal{U}$ and $\mathcal{V}$ are nonprincipal.6.1) implies that, under $\mathrm{CH}$, ${\mathcal{M}}_{\mathcal{U}}$ has $2^{\aleph_{1}}$ automorphisms and that ${\mathcal{M}}_{\mathcal{U}}$ and ${\mathcal{N}}_{\mathcal{V}}$ are isomorphic if and only if ${\mathcal{M}}$ and ${\mathcal{N}}$ are elementarily equivalent. Whether this conclusion is true without the $\mathrm{CH}$ depends on whether the theory of ${\mathcal{M}}$ is stable or not (see [69] and [73]).
@@ 11.0.9. Maximal rigidity
In Example 2.8 we saw that by [168] another quotient group, $S_{\infty}/H$, (where $H$ denotes the normal subgroup of all finitely supported permutations of ${\mathbb{N}}$) has the property that (provably in $\mathrm{ZFC}$) each of its automorphisms is given as conjugation by an almost permutation, and in particular it has a continuous lifting. One consequence of this strong rigidity result is the following. Fix a pair $f$, $g$ in $S_{\infty}$, and let ${\mathcal{M}}=(S_{\infty},f)$ and ${\mathcal{N}}=(S_{\infty},g)$. Then ${\mathcal{M}}/H$ and ${\mathcal{N}}/H$ are isomorphic if and only if $f$ and $g$ are conjugate (by an almost permutation) in the quotient, and in particular the question of the isomorphism of these structures is independent from $\mathrm{ZFC}$.
@@
What other structures (or categories) have this rigid behaviour, and is it possible to give a general characterisation of such structures?
@@ 12.1.2. BDF
The Brown–Douglas–Filmore question (§4.3), whether there exists an automorphism of the Calkin algebra that sends the unilateral shift to its adjoint, can be considered within the framework of (12.1). Considering ${\mathcal{M}}=(\mathcal{B}(H),s)$ and ${\mathcal{N}}=(\mathcal{B}(H),s^{*})$ as structures in the language of $\mathrm{C}^{*}$-algebras expanded by a constant interpreted by the shift or by its adjoint, the question is whether ${\mathcal{M}}/\mathcal{K}(H)$ is isomorphic to ${\mathcal{N}}/\mathcal{K}(H)$. The lamentable status of this question can be summarised in two remarks. First, it is not even known whether these two structures are elementarily equivalent. Second, even if they were, the Calkin algebra—and therefore each of these two structures—fails even to be countably homogeneous (see §6.3). Nevertheless, Woodin’s $\Sigma^{2}_{1}$ absoluteness theorem suggests that assuming the Continuum Hypothesis may help. In addition, forcing $\mathrm{CH}$ without adding reals preserves $\Sigma^{2}_{1}$ statements and therefore assuming $\mathrm{CH}$ does no harm in this context.
@@ 12.1.3. Non-isomorphism
In order to avoid metamathematical detours, instead of question (12.2) we consider a question that is, given the current state of the art in set theory, easier to analyse: When does /≇${\mathcal{M}}/{\mathcal{I}}\not\cong{\mathcal{M}}/{\mathcal{J}}$ hold in a forcing extension?